OpenStudy (anonymous):

Evaluate the following. Write your answer in scientific notation. (3 x 10^7)(6.1 x 10^3) A. 18.3 x 10^10 B. 18.3 x 10^21 C. 1.83 x 10^10 D. 1.83 x 10^11

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

\[3\times 10^7\times 6.1\times 10^3=3\times 6.1\times 10^{7+3}=18.3\times 10^{10}=1.83\times 10^{11}\]

OpenStudy (anonymous):

thnk u.